OpenStudy (anonymous):

How many solutions does the equation 5x + 3x –4 = 10 have? Zero One Two Infinitely many

OpenStudy (solomonzelman):

if you can't tell me right away, then lets solve this equatin.

OpenStudy (solomonzelman):

Do you see any like terms on the left hand side of the equation ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

I will color the like terms in blue \(\large\color{black}{ \displaystyle \color{blue}{5x} + \color{blue}{3x} -4 = 10 }\)

OpenStudy (solomonzelman):

Add the x 's. What is \(\normalsize\color{black}{ \displaystyle \color{blue}{5x+} \color{blue}{3x} }\) ?

OpenStudy (anonymous):

8x

OpenStudy (solomonzelman):

yes, so \(\large\color{black}{ \displaystyle \color{blue}{5x} + \color{blue}{3x} -4 = 10 }\) \(\large\color{black}{ \displaystyle \color{blue}{8x} -4 = 10 }\)

OpenStudy (solomonzelman):

now, you have variables (x's ) and numbers (the -4 in this case) on left side and on right side you have a number. can you eliminate the variables for me?

OpenStudy (solomonzelman):

lets add 4 to both sides \(\large\color{black}{ \displaystyle 8x-4 \color{darkgoldenrod}{\bf +4} = 10\color{darkgoldenrod}{\bf +4} }\)

OpenStudy (solomonzelman):

what is -4+4 = ?

OpenStudy (anonymous):

0

OpenStudy (solomonzelman):

yes, and 10+4

OpenStudy (solomonzelman):

is?

OpenStudy (anonymous):

14

OpenStudy (solomonzelman):

Yup !! \(\large\color{black}{ \displaystyle 8x-4 \color{darkgoldenrod}{\bf +4} = 10\color{darkgoldenrod}{\bf +4} }\) \(\large\color{black}{ \displaystyle 8x+0= 14 }\) so we now have just \(\large\color{black}{ \displaystyle 8 \times x = 14 }\)

OpenStudy (solomonzelman):

by what do you divide on both sides, to solve for x (to isolate the x) ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so is there only one posibiltity to make an equation?

OpenStudy (anonymous):

or is there more?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 8x= 14 }\) alright, I will tell you. We have to divide by 8 on both sides \(\large\color{black}{ \displaystyle 8x \color{magenta}{\div 8}= 14 \color{magenta}{\div 8} }\)

OpenStudy (solomonzelman):

yes, correct there is only one solution

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 8x \color{magenta}{\div 8}= 14 \color{magenta}{\div 8} }\) \(\large\color{black}{ \displaystyle \cancel{8}x \color{magenta}{\div \cancel{8}}= 14 \color{magenta}{\div 8} }\) \(\large\color{black}{ \displaystyle x= 14 \color{magenta}{\div 8} }\)

OpenStudy (solomonzelman):

x=14/8 x=7/4

OpenStudy (solomonzelman):

and it is x=7/4

OpenStudy (anonymous):

ok thanks for the help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!SolomonZelman

OpenStudy (solomonzelman):

sure